Calculate the expected vapor pressure at 25 °C for a solution prepared by dissolving 235 g ethanol (M = 46 g/mol) in 500 mL of water. At 25 °C, the density of water is 0.9971 g/cm3 and the vapor pressure is 23.76 torr.

3.7 torr

20.06 torr

10.6 torr

12 torr
Answer #1

Using Raoult's law, P​​​​​​soln = X​​​​​​H2O P​​​​​​H2O

to find the mole fraction of H2O, we have to find the number of moles .

Moles of ethanol = 235/46 = 5.108 mol

Mass of water = volume*density = 500*0.9971 = 498.55 g

Moles of water= 498.55/18 = 27.697 mol

X​​​​​​H2O = 27.697/(27.697+5.108) = 0.844

P = 0.844*23.76 = 20.06 torr

Answer:20.06 torr